Biphenyl, C12H10, is a nonvolatile, nonionizing solute that is soluble in benzene, C6H6. at 25 degrees Celsius, the vapor pressure of pure benzene is 100.84 torr. what is the vapor pressure of a solution made from dissolving 10.1 g of biphenyl in 28.5 g of benzene?
Answer #1
Concept and reason

The concept used is to calculate the vapor pressure of the solution.

Fundamentals

A solute is a substance that is dissolved in another substance. It is present in smaller quantities.

A solvent is a liquid in which the solute is dissolved. It is present in larger amounts.

According to Raoult’s law, when a non-volatile solute is mixed in a solvent, the vapor pressure of a solution is equal to the vapor pressure of the pure solvent multiplied by its mole fraction.

It is expressed as follows:

Psolution=xsolvent×Psolvent{P_{{\rm{solution}}}} = {x_{{\rm{solvent}}}} \times {P_{{\rm{solvent}}}}

Moles fraction of a solvent is the number of moles of solvent divided by the total moles of solute and solvent.

It is expressed as follows:

xsolvent=nsolventnsolvent+nsolute{x_{{\rm{solvent}}}} = \frac{{{n_{{\rm{solvent}}}}}}{{{n_{{\rm{solvent}}}} + {n_{{\rm{solute}}}}}}

Here in this question, benzene is the solvent and biphenyl is the solute. So, the vapor pressure of the solution is as follows:

Psolution=xbenzene×Pbenzene{P_{{\rm{solution}}}} = {x_{{\rm{benzene}}}} \times {P_{{\rm{benzene}}}}

1.

Mass of biphenyl = 10.1g10.1{\rm{ g}}

Molar mass of biphenyl = 154g/mol154{\rm{ g/mol}}

Molesofbiphenyl=10.1g154g/mol=0.066mol\begin{array}{l}\\{\rm{Moles of biphenyl}} = \frac{{10.1{\rm{ g}}}}{{154{\rm{ g/mol}}}}\\\\{\rm{ = 0}}{\rm{.066 mol}}\\\end{array}

Mass of benzene = 28.5g28.5{\rm{ g}}

Molar mass of benzene = 78g/mol78{\rm{ g/mol}}

Molesofbenzene=28.5g78g/mol=0.37mol\begin{array}{l}\\{\rm{Moles of benzene}} = \frac{{28.5{\rm{ g}}}}{{78{\rm{ g/mol}}}}\\\\{\rm{ = 0}}{\rm{.37 mol}}\\\end{array}

The mole fraction of benzene is calculated as follows:

Molefractionofbenzene(xbenzene)=0.37mol(0.37+0.066)mol=0.85\begin{array}{l}\\{\rm{Mole fraction of benzene}}\left( {{x_{{\rm{benzene}}}}} \right) = \frac{{0.37{\rm{ mol}}}}{{\left( {0.37 + 0.066} \right){\rm{ mol}}}}\\\\{\rm{ = 0}}{\rm{.85}}\\\end{array}

Vapor pressure of pure benzene = 100.84torr100.84{\rm{ torr}}

The vapor pressure of the solution is calculated as follows:

Psolution=(0.85)×(100.84torr)=85.71torr\begin{array}{l}\\{P_{{\rm{solution}}}} = \left( {0.85} \right) \times \left( {100.84{\rm{ torr}}} \right)\\\\{\rm{ = 85}}{\rm{.71 torr}}\\\end{array}

Ans: Part 1

Therefore, the vapor pressure of the solution is 85.71torr85.71{\rm{ torr}} .